Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-2021-05-10

Observers: Daniele Bjørn Malesani, Giorgos Leloudas, Miika Pursiainen
Support: Miika Pursiainen, Panos Charalampopoulos
TAT: Giuliano Pignata, Santiago Gonzalez, Jacob Teffs

EFOSC Calibrations: 
Bias, dome_flats, arcs. - however there were problems with the slit alignment (see below). So we will repeat calibration in the morning to be in the safe side (maybe arcs are affected in the evening set)
Bias, dome_flats, arcs. - calibrations repeated in the morning


Weather:
Start: Humidity ~30 %, wind < 5 m/s
- cloudy!
Middle:
Humidity 25 % , wind 8-10 m/s , clear
End: Humidity 18 % , wind 8-10 m/s , clear

Comments:
The first part of the night was affected by several problems:
- there were thick clouds in twilight - luckily they went away;
- we started with EFOSC2; the images of the standard star were highly out of focus;
- so the astronomer attempted the (longer) operation to focus the telescope;
- while doing this, the system failed, it's apparently an ongoing issue of connectivity going on for several days;
- so they had to restart the system; focus went finally well, PSF 1.2";
- this is when things went really bad: the slit turned out to be misaligned; after acquiring, no light from the object would come in the slit;
- we tried a couple of OBs, until they rebooted totally the system again, but to no avail;
- so we finally gave up and switched to SofI; we lost overall the first 2.5 hr of the night;
- later, the telescope operator figured out what was the problem, and the slits were aligned
- SOFI observations went fine.
- after completing the SOFI observations, we succesfully switched back to EFOSC2. Operations proceeded but with several problems causing loss of time:
- flats after g16 observations had to be repeated manually as the calibration units were not properly in place (apparently related to the guide probe problem). The TiO was doing this manually but there was problem with the preset.
- after target 2021fxf we aborted the g16 flats as they were taking disproportionately too much time and skipped to the next target.
- there were continuous problems with the guide probe. The TiO made several attempts but apparently one unit within the guide probe is stuck. We proceeded with tracking but not guiding. This seemed to work fine.

 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
EFOSC2
        
23:17
Telescope opens






Cloudy until now
23:17L745a Standardg13, g11, g16 1", 1.5"Various
1.2
 Attempt to focus on the standard stars.
Numerous problems arise, see above
         
SofI
        
01:47
SN2021bnw
Follow-upJHK---
1500
1.5
?

02:21
SN2021hiz
Follow-up JH
---
1x900s1.7
2.0

02:27
SN2018ubFollow-upJHK---
1x2400s1.5
1.7

02:56 SN2021kut Follow-up JH --- 760
  1.5  (DIMM)
 
03:13 SN2021jap Follow-up JH --- 760  1.5  (DIMM) 
03:31 SN2021jvs Follow-up JH --- 400  1.5  (DIMM) 
03:45 SN2021ipb     Follow-up JH --- 760  1.5  (DIMM) 
04:00 SN2021hem Follow-up JH --- 760  1.5  (DIMM) 
04:17 SN2021ixi Follow-up JH  --- 760  1.2  (DIMM) 
04:33 SN2021koq Follow-up
 JH  --- 200  1.2  (DIMM) 
         
 EFOSC2        
 05:05 AT2017des  Follow-up  V
   --- 60 s
  1.2  (DIMM) Only executed acquisition to inspect, not spectrum
 05:10 LTT379 Standard g13, g11, g16 1", 1.5" Various  1.2  (DIMM) 
 05:30 AT2021lmt Classification g13 1" 600 s
 1.1 1.4 
 05:44 SN2021fxy         Follow-up g11, g16 1"
 2 x 900 s
 1.4 1.4 Needed to repeat g16 calibration lamps.
 06:40 SN2021fxf Follow-up g11, g16 1" 2 x 1800
 1.2 1.4Problems with finding guiding star. After some time it worked but we are using tracking mode (not guiding).
Again problem with calibration lamps. Aborting those, taking too much time
 08:30     AT2020zso Follow-up g13 1" 2700 s
 1.5 1.5 (DIMM)
 Again problems with the guide probe. Losing time. Tracking not guiding
 09:50 AT2021lwx  Classification g13 1" 600 s
 1.9 1.5 (DIMM) 
 10:05 EG131 Standard  g13, g11, g16 1", 1.5" Various 1.1 1.5 (DIMM) 
         
         
Comments